A company that produces and markets video game systems wishes to assess its customers' level of satisfaction with a relatively new model, the XYZ-Box. In the six months since the introduction of the model, the company has received 73,219 warranty registrations from purchasers. The company will select a random sample of 65 of these registrations and will conduct telephone interviews with the purchasers. Specifically, each purchaser will be asked to state his or her level of agreement with each of the seven statements listed on the survey instrument given in the following table.. Here, the level of agreement for each statement is measured on a 7-point Likert scale Purchaser satisfaction will be measured by adding the purchaser's responses to the seven statements. It follows that for each consumer the minimum composite score possible is 7 and the maximum is 49. Furthermore, experience has shown that a purchaser of a video game system is "very satisfied" if his or her composite score is at least 42.


The Video Game Satisfaction Survey Instrument

StatementStrongly Disagree




Strongly Agree
The game console of the XYZ-Box is well designed.1234567
The game controller of the XYZ-Box is easy to handle.1234567
The XYZ-Box has high quality graphics capabilities.1234567
The XYZ-Box has high quality audio capabilities.1234567
The XYZ-Box serves as a complete entertainment center.1234567
There is a large selection of XYZ-Box games to choose from.1234567
I am totally satisfied with my XYZ-Box game system.1234567

Suppose that when the 65 customers are interviewed, their composite scores are as given in the following table.

Composite Scores for the Video Game Satisfaction Rating Case
3845434039
4648454344
3839474341
4441414645
3838494339
48464839
45404438
46404338
43414441
38404538
43404238
46473938
48424446
46454743
45403846

Using the data, estimate limits between which most of the 73,219 composite scores would fall. Also, estimate the proportion of the 73,219 composite scores that would be at least 42. (Round your proportion of scores answer to 2 decimal places.)


Most of the scores would fall between _______ and _______  

An estimate of the proportion of scores that would be at least 42 = _______ %
Answer #1

Answer:

score

n

65

mean

42.63

sample standard deviation s

3.33

Most of the scored falls between 3 standard deviation limits.

   Mean – 3s

32.63

   mean + 3s

52.63

Most of the scored would fall between 32.63 and 52.63

since 73219 is very large, we use z distribution to find proportion.

Z value for 42 is z =(42-42.63)/3.33 = -0.19

P( x ≥42) = P( z > -0.19) =0.5753

An estimate of proportion of scores that would be atleast 42 = 57.53%
Answer #2

so to do this question, the answer to the "most" part is just the lowest number and highest number. And then second part is literally counting the numbers less than 42 and dividing it by 65 than 1 minus that decimal.

 eg: 35 numbers are less than 42 therefore 35/65=x and then 1-x *100 is the answer

source: my brain
answered by: shordy speero